# Temporal reasoning for timeline summarisation in social media

### v1

@@ Abstract
This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarization, the task of summarising long texts containing sequences of events, particularly social media threads . We introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships. Our approach then combines temporal reasoning with timeline summarization through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarization. Experimental results demonstrate that our model achieves superior performance on mental health-related timeline summarization tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance of leveraging temporal reasoning to improve timeline summarisation.
@@ 1 Introduction
Timeline summarization organizes and presents a sequence of events in a coherent and concise manner Steen and Markert (2019); Li et al. (2021); Hu et al. (2024). It involves extracting event-related timelines and then summarising them Hu et al. (2024); Rajaby Faghihi et al. (2022). Researchers generally create event graphs Li et al. (2021) or cluster event related timelines Hu et al. (2024) to identify relevant events. Recent work Song et al. (2024) has introduced the challenging task of social media timeline summarisation, especially in the context of capturing fluctuations in individuals’ state of mind as reflected in posts shared online over time. In these posts, numerous events may occur without explicit timestamps, requiring contextual inference to determine their chronological sequence. Moreover, mental health-related events are not easy to identify: they can be connected to an individual’s emotions, interpersonal interactions, and the entire timeline is necessary to provide enough contextSong et al. (2024). It is particularly challenging to identify events pertaining to psychological states and to extract these from posts. When generating mental health related summaries from longitudinal posts, models need to understand related events and maintain temporal consistency to make inferences. This raises the question of whether temporal reasoning can be leveraged to enhance the quality of complex timeline summaries.
@@
Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a). (Zhou et al., 2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity). Subsequently, Tan et al. (2023a); Jain et al. (2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to the temporal commonsense aspects. LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al. (2023). Thus temporal commonsense reasoning is beneficial for timeline summarisation, as it helps maintain temporal consistency and the correct event order Wenzel and Jatowt (2023b); Vashishtha et al. (2020). Recent work Chan et al. (2024); Feng et al. (2023); Chen et al. (2024); Zhang et al. (2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how enhancing these abilities impacts downstream tasks, such as timeline summarisation.
@@
Here we propose combining temporal reasoningwith timeline summarisation using LLMs, to enhance the generation of timeline summaries. Specifically, we first fine-tune a teacher model using a novel temporal reasoning dataset (NarrativeReason) and then distill temporal reasoning knowledge into a smaller student model, which is simultaneously fine-tuned on the timeline summarisation task.
@@
    Based on the timelines created by NarrativeTime on the TimeBankNT corpus Rogers et al. (2024), we develop a new dataset NarrativeReason for temporal reasoning. Unlike existing temporal reasoning datasets Tan et al. (2023a); Chu et al. (2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than event pairs. This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
@@
    We apply the fine-tuned LLM to a completely different domain from what it is trained on, specifically generation of mental health related timeline summaries. Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al. (2024). Not only does it generate more accurate summaries but it also reduces hallucinations in LLMs.
@@
    We show why knowledge distillation works well, and how it induces better learned representations, through activation analysis of the fine-tuned LLM.
@@ 2 Related Work
Temporal reasoning for LLMs Temporal reasoning in Natural Language Processing (NLP) is the ability to understand and process information related to time within natural language text. It includes reasoning about the chronology and duration of events, and understanding and capturing different temporal relations Vashishtha et al. (2020). Despite the impressive performace of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g. translation, generation), they have been shown to perform suboptimally in temporal reasoning Wang and Zhao (2024); Chu et al. (2024); Qiu et al. (2023). However the ability to perform temporal reasoning is crucial for understanding narratives Nakhimovsky (1987); Jung et al. (2011a); Cheng et al. (2013), answering questions Bruce (1972); Khashabi (2019); Ning et al. (2020), and summarising events Jung et al. (2011b); Vashishtha et al. (2020). Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs. Xing and Tsang (2023) Huang et al. (2024). To increase understanding of temporal expressions, Tan et al. introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event). TEMPREASON was used to fine-tune a LLM to improve the model’s temporal reasoning, and they analysed the performance of different LLMs on this dataset showing it is challenging for LLMs to capture the temporal relationships between different events. Xiong et al. use the aligned timeline to help the LLM to understand temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge. The temporal relationship between events is inferred based on specific times (e.g., the year the events occurred). However in a narrative, events often occur without a clear indication of time.
Temporal reasoning for summarization Jung et al. developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied tomedical records, presenting medical history in a more intuitive way. They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis. Temporal reasoning can help the LM analyze temporal relationships between different events, which is very beneficial for event or news summarisation Vashishtha et al. (2020). However, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
@@ 3 Methodology
Figure 1: Overview of proposed method. (a) represents fine-tuning the teacher model on the temporal reasoning dataset.
@@ 3.1 Proposed architecture
(1) Improving temporal reasoning. We fine-tune a large LLM as a teacher model on the ‘NarrativeReason’ dataset §3.2.
@@
(2) After fine-tuning the teacher model, we freeze its parameters. At this stage, we fine-tune a student model (a smaller LLM) on the the task of timeline summarisation Chen et al. (2023a) from news. During this process, the teacher model transfers temporal reasoning knowledge to the student model, while the student simultaneously leverages the acquired temporal reasoning knowledge to perform timeline summarisation. For knowledge distillation (KD), we adopt three different strategies: Neuron Selectivity Transfer (NST), Contrastive Representation Distillation (CRD) and Probabilistic Knowledge Transfer (PRT). When generating the timeline summary, we conduct experiments on the TalkLife dataset, a very different domain to the one the student is trained on. We prompt the student model to generate mental health related summaries pertaining to aspects such as diagnostic states, inter- and intra- personal relationships and fluctuations in mood Song et al. (2024).
3.2 Teacher Model
@@
Here the goal is to improve an LLM’s temporal reasoning. There is evidence showing that fine-tuning on datasets such as TEMPLAMA may enable an LLM to memorise the most frequent answer rather develop temporal reasoning Tan et al. (2023b). In other words, the model does not truly understand temporal relationships, such as "before" and "after". Most temporal reasoning datasets involve pairs of events rather than multiple events. However, processing a sequence of events requires more intricate reasoning, including recognising patterns, dependencies, and causal chains among multiple events. This is useful for more sophisticated tasks such as narrative comprehension and timeline summarisation, where understanding the full sequence of events is crucial.
To prevent the LLM from learning shortcuts and memorising the most frequent answer, we created a temporal reasoning dataset NarrativeReason, which contains relationships between a series of events based on a given story.
@@
Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al. (2024), which re-annotated TimeBankDense Cassidy et al. (2014) with a timeline-based annotation framework, NarrativeTime. They annotated all possible temporal links (TLINKS) and the temporal relationships of all events occurring within astory. They thus provide a complete temporal relation for the sequence of events in the text, rather than just the temporal relation between event pairs. That means they provide a clear timeline for these events. However, they only used one word (e.g. "fallen") to represent an event, which does not fully capture the complete picture of the event (e.g. "the Indonesian stock market has fallen by twelve percent"). In general, an event is considered an action involving its corresponding participants, and is marked by annotating a representative expression of the event. In the context of temporal reasoning, events are usually the head of the verb phrase. Ning et al. (2018); Pustejovsky et al. (2003).
@@
Therefore, we filtered the dataset TimeBankNT, keeping only annotated verbs to represent events. In order To represent a complete event, we use these verbs as triggers to extract relational triplets e.g. <Indonesian stock market value, fallen, by twelve percent> to represent the event fallen. Then, we use these triplets to construct the temporal relationship of a series event. (e.g. Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
@@
Dataset construction In a story, we consider the temporal relations for all events, and construct event-event relation question/answer pairs, which addrees the chronological relation between events, such as ‘before’, ‘after’, ‘during’, and ‘simultaneous’ Tan et al. (2023a). Specifically, we obtain the temporal relations of all events and then use questioanswering prompts to reconstruct the dataset. Question: your task is to identify the temporal relation between EVENT A and EVENT B: based on the Story: STORY. Answer: EVENT A temporal relation (BEFORE/ AFTER/ INCLUDES/ IS_INCLUDED/ SIMULTANEOUS) EVENT B. Although a single question-answer pair is used to determine the temporal relationship between a pair of events, for a complete story, we construct multiple question-answer pairs to cover the temporal relationsamong all events. This ensures that the model can clearly outline the temporal relations between all events in the story.
@@
Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al. (2022). The input and output of the model are the temporal questions and corresponding answers. Our experiments reveal that fine-tuning on this dataset brings improvements to other temporal reasoning tasks (Appendix A.2) .
Figure 2: The temporal relationship between events. We represent events triggered by verbs using relational triplets, as shown in the middle column. In the right column, we list a series of events and use arrows of different colors to represent the temporal relations between these events. In this example, we use blue, green and purple arrows to indicate ’Simultaneous’, ’After’ and ’Before’ respectively.
@@ 3.3 Student Model
After we fine-tune a teacher model on the reconstructed dataset, we transfer the temporal reasoning knowledge to a student model, while, also fine-tuning the student on a news timeline summarisation dataset Chen et al. (2023b). Thus we aim for the student to learn temporal reasoning and use this ability to generate timeline summaries. We fine-tune Phi-3-mini-4k-instruct as a student model. We use three knowledge distillation (KD) objectives to transfer knowledge from the teacher to the student: Neuron Selectivity Transfer (NST) Huang and Wang (2017), transfers heatmap like spatial activation patterns of teacher neurons to student neurons; Contrastive Representation Distillation (CRD) Tian et al. (2019), maximises the mutual information between the teacher and student representations with contrastive learning; Probabilistic Knowledge Transfer (PRT) Passalis and Tefas (2018), matches the probability distribution of the data in the feature space.
@@
The core idea is to match the probability distribution of the data in the feature space between teacher and student models. However, learning a significantly smaller model that accurately recreates the whole geometry of a complex teacher model is often impossible. Passalis and Tefas (2018) uses the conditional probability distribution to describe the samples. Here, $\mathbf{Y_{t}}={\{\mathbf{y_{t}}}_{1},{\mathbf{y_{t}}}_{2},...,{\mathbf{y_{t}}}_{l}\}\in\mathbb{R}^{vocab_{t}}$ denote the output logits of the teacher model, and $\mathbf{Y_{s}}={\{\mathbf{y_{s}}}_{1},{\mathbf{y_{s}}}_{2},...,{\mathbf{y_{s}}}_{l}\}\in\mathbb{R}^{vocab_{s}}$ denote the output logits of the student model, where ${\mathbf{y_{t}}}$/${\mathbf{y_{s}}}$ is vector and $l$ is the length of sentences, ${vocab_{t}}$ and ${vocab_{t}}$ are the vocabulary sizes of teacher and student models respectively. We can define the conditional probability distribution for the teacher model as Eq. 1, and student model as Eq. 2.
@@
${q_{i|j}}=\frac{K(\mathbf{x_{s}}_{i},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{x_{s}}_{k},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}}$ (2)
In these two equations (Eq. 1 and Eq. 2), we use the cosine similarity as a kernel metric allows for more robust affinity estimations.
${K_{consine}({\mathbf{y_{t}}_{i}},{\mathbf{y_{t}}_{j}})}=\frac{1}{2}(\frac{{\mathbf{y_{t}}_{i}^{\mathrm{T}}}\mathbf{y_{t}}_{j}}{\|{\mathbf{y_{t}}_{i}}\|{{}_{2}}\|{\mathbf{y_{t}}_{j}}\|{{}_{2}}}+1)\in[0,1].$
We use Kullback-Leibler (KL) divergence to calculate the distance between the conditional probability distributions of the teacher model and the student model, as follows:
@@
matches the distributions of neuron selectivity patterns between teacher and student networks. We transfer the last hidden layer $\mathbf{T}=\mathbf{t(x)}$ of the teacher model to the last hidden layer $\mathbf{S}=\mathbf{s(x)}$ of the student model given input text $\mathbf{x}$. Specifically, we transfer neuron selectivity knowledge from $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$ to $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$, where $N$ and $M$ are the hidden state dimensions. Then we follow the idea in Huang and Wang (2017) using Maximum Mean Discrepancy (MMD) to calculate the distance between these two distributions. The basic idea of MMD is that if the moments of two random variables are the same for all orders, then the two distributions are identical $L_{MMD^{2}}(\mathbf{t},\mathbf{s})$.
@@
$-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}k\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}}.$
In this formula we use a Gaussian Kernel ${k(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
@@
maximises the lower-bound to the mutual information between the teacher and student representations. Here, we follow Tang et al. (2021) to sample one positive pair from the joint distribution $p(\mathbf{S},\mathbf{T})=q(\mathbf{S},\mathbf{T}|positive)$ for $N$ negative pairs sampled from the product of marginal $p(\mathbf{S})p(\mathbf{T})=q(\mathbf{S},\mathbf{T}|negtive)$, where $N$ is the batch size. We can maximize the lower bound of mutual information by minimizing the following loss function:
@@
In Eq 3, function $h$ should satisfy $h:\{\mathbf{s},\mathbf{t}\}\rightarrow[0,1]$,
@@
where M is the cardinality of the dataset, and we need to normalize $\mathbf{s}$ and $\mathbf{t}$ by $L$-2 norm before the inner product.
The knowledge distillation method transfers temporal reasoning knowledge from the teacher model to the student model. At the same time, we aim for this knowledge to benefit the timeline summarization task. Therefore, we fine-tune the student model on the timeline summarization dataset §4.1, enabling it to simultaneously learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarization information.

### v2

@@ Abstract
This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarisation, the task of summarising long texts containing sequences of events, such as social media threads. We first introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships. Our approach then combines temporal reasoning with timeline summarisation through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarisation. Experimental results demonstrate that our model achieves superior performance on out-of-domain mental health-related timeline summarisation tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance and generalisability of leveraging temporal reasoning to improve timeline summarisation.
@@ 1 Introduction
Timeline summarisation organizes and presents a sequence of events in a coherent and concise manner Steen and Markert (2019); Li et al. (2021); Hu et al. (2024). It involves extracting event-related timelines and then summarising them Hu et al. (2024); Rajaby Faghihi et al. (2022). Researchers generally create event graphs Li et al. (2021) or cluster event related timelines Hu et al. (2024) to identify relevant events. Recent work Song et al. (2024) has introduced the challenging task of social media timeline summarisation, especially in the context of capturing fluctuations in individuals’ state of mind as reflected in posts shared online over time. In these posts, numerous events may occur without explicit timestamps, requiring contextual inference to determine their chronological sequence. Moreover, mental health-related events are not easy to identify: they can be connected to an individual’s emotions, interpersonal interactions, and the entire timeline is necessary to provide enough context Song et al. (2024). It is particularly challenging to identify events pertaining to psychological states and to extract these from posts. When generating mental health related summaries from longitudinal posts, models need to understand related events and maintain temporal consistency to make inferences. This raises the question of whether temporal reasoning can be leveraged to enhance the quality of complex timeline summaries.
@@
Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a). Zhou et al. (2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity). Subsequently, Tan et al. (2023); Jain et al. (2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to temporal commonsense aspects. LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al. (2023). Thus temporal commonsense reasoning is beneficial for timeline summarisation, as it helps maintain temporal consistency and the correct event order Wenzel and Jatowt (2023b); Vashishtha et al. (2020). Despite the evidenced connection between temporal reasoning and timeline summarisation, recent work Chan et al. (2024); Feng et al. (2023); Chen et al. (2024); Zhang et al. (2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how they impact downstream tasks, such as timeline summarisation.
@@
Here we propose combining temporal reasoning with timeline summarisation using LLMs, to enhance the generation of timeline summaries. Specifically, we first fine-tune a teacher model using a novel temporal reasoning dataset (NarrativeReason) and then distill temporal reasoning knowledge into a smaller student model, which is simultaneously fine-tuned on the timeline summarisation task, trained and tested in separate domains.
@@
    Based on the timelines derived in NarrativeTime Rogers et al. (2024) from the TimeBankNT corpus Cassidy et al. (2014), we develop a new dataset for temporal reasoning, NarrativeReason. Unlike existing temporal reasoning datasets Tan et al. (2023); Chu et al. (2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than distinct event pairs. This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
@@
    We fine-tune a large LLM on NarrativeReason and distill its knowledge to a smaller model, which is fine-tuned for the task of timeline summarisation. The resulting fine-tuned smaller LLM is applied to a completely different domain from the one it is trained on. Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al. (2024). Not only does it generate more accurate summaries but it also reduces hallucinations in LLMs.
@@
    We show why knowledge distillation works well, and how it induces better learned representations, through activation analysis of the fine-tuned smaller LLM.
@@ 2 Related Work
Temporal reasoning for LLMs Temporal reasoning in Natural Language Processing (NLP) is the ability to understand and process information related to time within natural language text. It includes reasoning about the chronology and duration of events, and understanding and capturing different temporal relations Vashishtha et al. (2020). Despite the impressive performance of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g. translation, generation), they have been shown to perform sub-optimally in temporal reasoning Wang and Zhao (2024); Chu et al. (2024); Qiu et al. (2023). However the ability to perform temporal reasoning is crucial for understanding narratives Nakhimovsky (1987); Jung et al. (2011a); Cheng et al. (2013), answering questions Bruce (1972); Khashabi (2019); Ning et al. (2020), and summarising events Jung et al. (2011b); Vashishtha et al. (2020). Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs Xing and Tsang (2023) Huang et al. (2024). To increase understanding of temporal expressions, Tan et al. (2023) introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event). TEMPREASON was used to fine-tune a LLM to improve its temporal reasoning, and performance of different LLMs on this dataset showed it is challenging for LLMs to capture the temporal relations between different events. Xiong et al. (2024) use an aligned timeline to improve an LLM’s temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge. The temporal relationship between events is inferred based on specific times (e.g., year of event). However in a narrative, events often occur without a clear indication of time.
Temporal reasoning for summarisation Jung et al. (2011b) developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied to medical records, presenting medical history in a more intuitive way. They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relationships between different events, which is beneficial for event or news summarisation Vashishtha et al. (2020). However, despite the evidenced connection, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
@@ 3 Methodology
Figure 1: Overview of proposed method. (a) represents fine-tuning the teacher model on the temporal reasoning dataset; (b) In the KD process, we input the timeline summarisation dataset into both the Teacher and Student models, transferring temporal reasoning knowledge while using it to assist the timeline summarisation task to fine-tune Student model.
@@ 3.1 Proposed architecture
(1) Improving temporal reasoning. We fine-tune a large LLM as a teacher model on our ‘NarrativeReason’ dataset §3.2.
@@
(2) After fine-tuning the teacher model, we freeze its parameters. At this stage, we fine-tune a student model (a smaller LLM) on timeline summarisation from news Chen et al. (2023). During this process, the teacher model transfers temporal reasoning knowledge to the student model, while the student simultaneously leverages the acquired temporal reasoning knowledge to perform timeline summarisation. For knowledge distillation (KD), we adopt three different strategies: Neuron Selectivity Transfer (NST), Contrastive Representation Distillation (CRD) and Probabilistic Knowledge Transfer (PRT). When generating the timeline summary, we conduct experiments on the TalkLife dataset, which pertains to a very different domain (mental health) to the one the student is trained on (news). We prompt the student model to generate mental health related summaries pertaining to aspects such as diagnostic states, inter- and intra- personal relationships and fluctuations in mood following Song et al. (2024).
3.2 Teacher Model and NarrativeReason dataset
@@
Here the goal is to improve an LLM’s temporal reasoning. Evidence has shown that fine-tuning on datasets such as TEMPLAMA may enable an LLM to memorise the most frequent answer rather than develop temporal reasoning Tan et al. (2023). In other words, the model does not truly learn the meaning of temporal relations, such as "before" and "after". We hypothesise that this is because most temporal reasoning datasets involve pairs of events rather than multiple events. Processing a sequence of events requires more intricate reasoning, including recognising patterns, dependencies, and causal chains among multiple events. This is useful for more sophisticated tasks such as narrative comprehension and timeline summarisation, where understanding the full sequence of events is crucial. To prevent the LLM from learning shortcuts and memorising the most frequent answer, we created a temporal reasoning dataset NarrativeReason, which contains relations between a series of events based on a given narrative.
@@
Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al. (2024), which in turn had re-annotated TimeBankDense Cassidy et al. (2014) with a timeline-based annotation framework, NarrativeTime framework. Rogers et al. (2024) had annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get a NarrativeTime dataset. Here, TLINKs usually denote the event order information (e.g., before, after, during). Events more broadly and especially in the context of temporal reasoning, are represented as relation triples where the event trigger is the head of the verb phrase Ning et al. (2018); Pustejovsky et al. (2003), linking the corresponding arguments. However, NarrativeTime dataset  Rogers et al. (2024) only denotes the type of the event, annotated at the level of the verb, e.g. for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in NarrativeTime datasets as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent". However, this only constitutes a denser temporal annotation without providing event triples and is therefore unsuited for temporal reasoning training tasks.
@@
Thus we reconstructed NarrativeReason, to augment it with event triples that can be used for temporal reasoning training. Specifically, we filtered the NarrativeTime dataset, keeping only annotated verbs to represent events. In order to represent an event, we use verbs as triggers to extract relational triples e.g. <Indonesian stock market value, fallen, by twelve percent> to represent the event annotated in NarrativeTime dataset as fallen. Then, we use these triples to construct the temporal relations of a series event. (e.g. Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
@@
Dataset construction For a given narrative, we consider the temporal relations between all events, and construct question/answer pairs for event-event relations, addressing the chronological order of events (‘before’, ‘after’, ‘during’, and ‘simultaneous’ Tan et al. (2023)). Specifically, we obtain the temporal relations of all events and then use question answering prompts to reconstruct the dataset. Question: your task is to identify the temporal relation between EVENT A and EVENT B: based on the Story: STORY. Answer: EVENT A temporal relation (BEFORE/ AFTER/ INCLUDES/ IS_INCLUDED/ SIMULTANEOUS) EVENT B Tan et al. (2023). Although a single question-answer pair is used to determine the temporal relation between a pair of events, for a complete narrative, we construct multiple question-answer pairs to cover the temporal relations among all events. This ensures that the model is exposed to all temporal relations between all events in the story. Fig. 2 shows the data format of NarrativeReason.
@@
Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al. (2022). The input and output of the model are the temporal questions and corresponding answers. Our experiments show that indeed fine-tuning on NarrativeReason improves performance on established temporal reasoning tasks (Appendix A.2) .
Figure 2: The temporal relations between events. The text in the left column comes from NarrativeTime dataset, and we filtered it to keep only verbs. We represent events triggered by verbs using relational triples, as shown in the middle column. In the right column, we construct question/answer pairs for all events.
@@ 3.3 Student Model
After we fine-tune a teacher model on the NarrativeReason dataset, we transfer the temporal reasoning knowledge to a student model, while, also fine-tuning the student on a news timeline summarisation dataset Chen et al. (2023). Thus we aim for the student to learn temporal reasoning and also use this ability when generating timeline summaries (step (b) in Fig. 1). We fine-tune Phi-3-mini-4k-instruct as a student model. We use three knowledge distillation (KD) objectives to transfer knowledge from the teacher to the student: Neuron Selectivity Transfer (NST) Huang and Wang (2017), transfers heatmap like spatial activation patterns of teacher neurons to student neurons; Contrastive Representation Distillation (CRD) Tian et al. (2019), maximises the mutual information between the teacher and student representations with contrastive learning; Probabilistic Knowledge Transfer (PRT) Passalis and Tefas (2018), matches the probability distribution of the data in the feature space of teacher and student models.
@@
Learning a significantly smaller model that accurately recreates the whole geometry of a complex teacher model is often impossible. Passalis and Tefas (2018) uses the conditional probability distribution to describe the samples. Here, $\mathbf{Y_{t}}={\{\mathbf{y_{t}}}_{1},{\mathbf{y_{t}}}_{2},...,{\mathbf{y_{t}}}_{l}\}\in\mathbb{R}^{vocab_{t}}$ denote the output logits of the teacher model, and $\mathbf{Y_{s}}={\{\mathbf{y_{s}}}_{1},{\mathbf{y_{s}}}_{2},...,{\mathbf{y_{s}}}_{l}\}\in\mathbb{R}^{vocab_{s}}$ denote the output logits of the student model, where ${\mathbf{y_{t}}}$ and ${\mathbf{y_{s}}}$ are vectors and $l$ is sentence length, ${vocab_{t}}$ and ${vocab_{s}}$ are the vocabulary sizes of teacher and student models respectively. We can define the conditional probability distribution for the teacher model as Eq. 1, and student model as Eq. 2, where $K$ is a symmetric kernel with scale parameter $\sigma$.
@@
${q_{i|j}}=\frac{K(\mathbf{y_{s}}_{i},\mathbf{y_{s}}_{j};2\sigma_{s}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{y_{s}}_{k},\mathbf{y_{s}}_{j};2\sigma_{s}^{2})}}$ (2)
In equations Eq. 1 and Eq. 2, we use the cosine similarity kernel (no $\sigma$), allowing for more robust affinity estimations:
${K_{cosine}({\mathbf{y_{t}}_{i}},{\mathbf{y_{t}}_{j}})}=\frac{1}{2}(\frac{{\mathbf{y_{t}}_{i}^{\mathrm{T}}}\mathbf{y_{t}}_{j}}{\|{\mathbf{y_{t}}_{i}}\|{{}_{2}}\|{\mathbf{y_{t}}_{j}}\|{{}_{2}}}+1)\in[0,1].$
We use Kullback-Leibler (KL) divergence to calculate the distance between the conditional probability distributions of the teacher and student models:
@@
matches the distributions of neuron selectivity patterns between teacher and student networks. We transfer the last hidden layer $\mathbf{T}=\mathbf{t(x)}$ of the teacher model to the last hidden layer $\mathbf{S}=\mathbf{s(x)}$ of the student model given input text $\mathbf{x}$. Specifically, we transfer neuron selectivity knowledge from $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$ to $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$, where $N$ and $M$ are the hidden state dimensions. Then we follow Huang and Wang (2017) in using Maximum Mean Discrepancy (MMD) to calculate the distance between the activation patterns of student $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$ and teacher neurons $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$. Here, we use squared MMD to calculate the diatance between $\mathbf{t}$ and $\mathbf{s}$.
@@
$-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}K\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}},$
where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
@@
maximizes the lower-bound to the mutual information between the teacher and student representations. In other words, we would like to push the representations $\mathbf{s(x}_{i})$ and $\mathbf{t(x}_{i})$ closer together, while pushing apart $\mathbf{s(x}_{i})$ and $\mathbf{t(x}_{j})$. Here, we follow the sampling process of Tang et al. (2021), providing 1 positive pair for every $N$ (batch size) negative pairs. The positive pair is sampled from the joint distribution $p(\mathbf{S},\mathbf{T})=q(\mathbf{S},\mathbf{T}|positive)$, and $N$ negative pairs are drawn from the product of marginals $p(\mathbf{S})p(\mathbf{T})=q(\mathbf{S},\mathbf{T}|negative)$, where $q$ is a distribution denoting whether the ($\mathbf{S}$,$\mathbf{T}$) pair is drawn from the positive or negative pairs. We can maximize the lower bound of mutual information by minimizing the following loss function:
@@
In Eq 3, $h$ should satisfy $h:\{\mathbf{s},\mathbf{t}\}\rightarrow[0,1]$,
@@
where M is the cardinality of the dataset, and we need to normalize $\mathbf{s}$ and $\mathbf{t}$ by $L$-2 norm before taking the inner product.
The knowledge distillation process transfers temporal reasoning knowledge from the teacher to the student model. At the same time, we want this knowledge to benefit the timeline summarisation task. Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarisation information.